A car was bought on installment basis with a monthly installment of P30,000 for 60 months. If interest is 12% compounded annually, calculate the cash price of the car. {solve correctly with calculation}



1
Expert's answer
2021-07-12T11:46:25-0400

I believe that the student meant to put a cash purchase price of $10,000, at 12% annual APR, for 60 months. Then, calculate the corresponding payment. If any of that information (assumption) isn't correct, please let me know.

To solve this for total payments, we utilize the Future Value calculation of:

FV = C0 *(1+i)n

This, of course, must be adjusted for having monthly payments, which gives us:

FV = 10,000 * (1+0.01)60

FV = 10,000 * 1.816696699

Rounding this to only 2 significant digits, leave a future value of: $18,166.97